OpenStudy (anonymous):

What is the resulting equation when the expression for y in the second equation is substituted into the first equation? 3x + y = 1 y = 6 - 4x

OpenStudy (anonymous):

@santisanchez substitute

OpenStudy (anonymous):

3x+6-4=1 3x=-6+4+1 3x=-1?

OpenStudy (anonymous):

wait i forgot the x

OpenStudy (anonymous):

yeah you did

OpenStudy (anonymous):

-x+6=1

OpenStudy (anonymous):

hmm hey do you have choices ?

OpenStudy (anonymous):

3x + (6 - 4x) = 1 3x + 6 - 4x = 1 3x - 4x + 6 = 1 -x + 6 = 1 -x = -5 (-1) -x = -5(-1) x = 5

OpenStudy (anonymous):

thank you!